OpenStudy (anonymous):

Triangle ABC is similar to triangle DEF and has a ratio of similarity of 1:5. If side AB measures 6 mm, how long is side DE? don"t need answer need formula i learned a formula so far of cross multiplying 6/? 3/2 but its not helping with this ..

OpenStudy (anonymous):

AB/DE=BC/EF=CA/FB=1/5

OpenStudy (perl):